# Unleashing the Power of Task-Specific Directions in Parameter Efficient Fine-tuning

### v3

@@ B.1 How Do Task-specific Directions Arise in LoRA?
However, only the first conclusion of LoRA already presents significant contradictions. For a specific task, TSDs should indeed remain consistent and not depend on learned $\Delta\mathbf{W}$. Moreover, revisiting the first and the third conclusions, we could deduce that the top singular directions of $\Delta\mathbf{W}$ are among the directions of $\mathbf{W}$. Nonetheless, in practice, the singular directions of $\Delta\mathbf{W}$ are unlikely to align exactly with any specific directions in $\mathbf{W}$. As shown in our paper, although LoRA highlights the importance of TSDs for downstream tasks, it struggles with a clear definition of TSDs which leads to contradictions, and does not effectively outline how they can be leveraged, leading to potential under-utilization in practical applications
For more details on what they did and why they drew the conclusions, we encourage readers to refer to their original paper Hu et al. (2021), specifically Section 7.
B.2 Why should TSD be Established on $\mathbf{W}$?
@@
Indeed, the most accurate direction for TSD should be the direction of $\Delta\mathbf{W}^{*}$. However, since we do not have access to $\Delta\mathbf{W}^{*}$ or $\mathbf{W}^{*}$ during fine-tuning, we only have information about $\mathbf{W}$. Therefore, we establish TSD within the coordinate system based on the directions of $\mathbf{W}$.
@@
B.3 Why Must TSD be in the Core Bases Rather than Other Global Bases?
Task-specific directions are defined as a subset of the core bases. However, why can’t they be defined as a set of bases from the global bases of $\mathbf{W}$ instead? Indeed, other global bases might also experience significant coordinate changes. However, the primary reason for not considering these bases is the difficulty in defining them.
@@ B.3 Why Must TSD be in the Core Bases Rather than Other Global Bases?
Firstly, if TSDs were defined based on the change rates, since the coordinate values of $\mathbf{W}$ on the other global bases are zero, even a minimal projection of $\Delta\mathbf{W}$ onto these global bases could result in an excessively high change rate, which would be unreasonable. Secondly, if TSDs were defined based on the absolute magnitude of coordinate changes, consider a scenario where a coordinate value in the core basis and a coordinate value in the global basis change by the same amount. If the original coordinate value in the core basis was significantly larger than the change, the change, even if substantial, could be viewed as a minor perturbation relative to the original value. Therefore, the same magnitude of change does not necessarily imply that the corresponding directions are equally important. For these reasons, we define TSDs based solely on the rate of change and restrict their definition to the core bases.
Appendix C Details on LoRA-Dash
C.1 Algorithm of LoRA-Dash
The algorithm of LoRA-Dash is shown in Fig. 6.
C.2 Matrix Operation of Dash Phase
Since in Sec. 4.2, we have adopted vector calculations for change rates for clear representation. However, to enhance efficiency, we will transition to matrix computations for determining change rates.
Let ${\mathbf{U}}=\begin{bmatrix}\mathbf{u}_{1}&\ldots&\mathbf{u}_{n}\end{bmatrix}\in\mathbb{R}^{n\times n}$ and ${\mathbf{V}}=\begin{bmatrix}\mathbf{v}_{1}&\ldots&\mathbf{v}_{m}\end{bmatrix}\in\mathbb{R}^{m\times m}$ represent the left and right singular vectors for a weight matrix $\mathbf{W}\in\mathbb{R}^{n\times m}$ ($n<m$), we have
${\mathbf{U}}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{V}=\begin{bmatrix}\mathbf{u}_{1}^{\mathsf{T}}\\ \vdots\\ \mathbf{u}_{n}^{\mathsf{T}}\end{bmatrix}\Delta\mathbf{W}\begin{bmatrix}\mathbf{v}_{1}&\cdots&\mathbf{v}_{m}\end{bmatrix}=\begin{bmatrix}\mathbf{u}_{1}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{1}&\ldots&\mathbf{u}_{1}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{m}\\ \vdots&\mathbf{u}_{i}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{j}&\vdots\\ \mathbf{u}_{n}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{1}&\ldots&\mathbf{u}_{n}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{m}\end{bmatrix}.$ (7)
Therefore, we can derive the change rates $\mathbf{\delta}=\begin{bmatrix}\delta_{1}&\ldots&\delta_{n}\end{bmatrix}^{\mathsf{T}}$, which are the diagonal elements of the matrix $\mathbf{U}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{V}$.
Following a similar process, we can rewrite Eq. (5) as
$\sum_{i=1}^{s}\Delta{\sigma}_{i}\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}=\bar{\mathbf{U}}\Delta\mathbf{\Sigma}\bar{\mathbf{V}}^{\mathsf{T}},$ (8)
where $\bar{\mathbf{U}}=\begin{bmatrix}\bar{\mathbf{u}}_{1}&\ldots&\bar{\mathbf{u}}_{s}\end{bmatrix}$, $\bar{\mathbf{V}}=\begin{bmatrix}\bar{\mathbf{v}}_{1}&\ldots&\bar{\mathbf{v}}_{s}\end{bmatrix}$ and $\Delta\mathbf{\Sigma}={\rm diag}(\begin{bmatrix}\Delta\sigma_{1}&\ldots&\Delta\sigma_{s}\end{bmatrix})$.
Appendix D Details on Experiments
D.1 Implementation Details
We compare LoRA-Dash mainly with LoRA. The ranks of LoRA-Dash, LoRA and other methods are varied among {2, 4, 8, 16, 32, 64}. All methods are implemented using the publicly available PyTorch Paszke et al. (2019) framework, and all experiments are conducted on NVIDIA A100 GPUs [^4] We have also test that most of the experiments can also be conducted on one consumer GPU resource such as NVIDIA RTX3090.. The hyper-parameter $t$ of LoRA-Dash is set to 100, and $s=8$. We fine-tune all the layers of different models in all the experiments, and the detailed experimental settings are shown in the following Secs. D.4-D.6.
D.2 Details on Fully Fine-tuning LLaMA in Section 3.3
To further explore the properties of TSDs, we fully fine-tune LLaMA-7B Touvron et al. (2023a) on commonsense reasoning tasks. For the commonsense reasoning tasks, which consist of 8 distinct sub-tasks, each defined with specific training and testing sets, we have adopted the approach in Hu et al. (2023). Following this method, we amalgamated the training datasets from all sub-tasks to construct a comprehensive final training dataset. More details can be found in Sec. D.4. The experiment was conducted on 8 A100 GPUs.
The change rates in Fig. 1 are derived from the weights obtained from the self-attention value projection in the 16th layer. However, our validation confirms that the results and conclusions drawn from the diagrams are consistent, regardless of which layer or specific module within the attention mechanism is analyzed. For the original change rate $\delta$ for a core basis, it is scaled as $\ln(\delta+1)/3$.
@@ D.1 Implementation Details
Additionally, it is worth noting that the phenomenon presented in Fig. 1 is not an isolated case. Similar behavior has been observed in other tasks and models. For instance, as shown in Fig., results obtained by training Qwen2.5-7B Team (2024) on the math reasoning task Hu et al. (2023) exhibit patterns that are fundamentally consistent with our observations in Fig. 1.
@@
D.3 Details on Fine-tuning LLaMA with LoRA in Section 3.4
Figure 8: We track the recall of TSDs every 100 training steps in the query, key and value layers of the LLaMA-7B during LoRA fine-tuning, analyzing how well the continuous updated $\Delta\mathbf{W}$ captures TSDs. The left and right figure settings are the same as those in Fig. 2.
We fine-tune LLaMA-7B using LoRA under three ranks, and for every 100 training steps we record the top 8 core directions of $\mathbf{W}$ which exhibit the highest change rates based on the projection of continuously updated $\Delta\mathbf{W}$ (i.e., the directions related to top 8 largest $\delta_{i}=\mathbf{u}_{i}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{i}/\sigma_{i}$, where $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ and $\sigma_{i}$ are the corresponding left/right singular vectors and values of $\mathbf{W}$). These directions serve as the “TSDs” predicted by LoRA. We denote these directions as “launched TSDs” (LTSDs) for differentiation. We then identify the top 4 or 16 TSDs with the highest change rates in Sec. 3.3.
We conduct two experimental setups to explore that:
@@
The purpose of the first experiment is more intuitive, aiming to confirm whether the TSDs with the top highest change rates can indeed be captured by $\Delta\mathbf{W}$ from LoRA. The second experiment, which may seem a bit more obscure, is designed to address concerns that, besides capturing the top-ranked TSDs, the LTSDs identified might include irrelevant directions. Therefore, this second setup allows us to examine whether the LTSDs predominantly fall within the upper echelons of TSDs based on their change rates, providing insight into the overall relevance and accuracy of the LTSDs identified by LoRA. Specifically, we employ two metrics to quantify the quality of LTSDs:
  • Accuracy: Measures how many of the 8 LTSDs are correctly included in the top 16 TSDs.
@@
The results presented in Figs 2 and 8 lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSDs:

### v4

@@ 2.4 Rethinking “Task-Specific Directions” in LoRA
However, even the first conclusion of LoRA presents significant contradictions. For a specific task, TSDs should indeed remain consistent and not depend on learned $\Delta\mathbf{W}$. Moreover, revisiting the first and the third conclusions, we could deduce that the top singular directions of $\Delta\mathbf{W}$ are among the directions of $\mathbf{W}$. Nonetheless, in practice, the singular directions of $\Delta\mathbf{W}$ are unlikely to align exactly with any specific directions in $\mathbf{W}$ [^1] For more details on what they did and why they drew the conclusions, we encourage readers to refer to their original paper [13], specifically Section 7..
Indeed, we suggest that “TSDs” are intuitive depictions of the low-dimensional manifold [18, 19]. Given that the most impactful transformations occur within the low-dimensional manifold during fine-tuning, TSDs are also significant in PEFT. However, the description of TSDs in LoRA presents several contradictions that contribute to confusion. This confusion complicates the identification of TSDs and the subsequent utilization of them.
We believe that these issues are mainly due to an unclear definition of TSDs. Consequently, to clearly define TSDs and enhance their practical application, we have discarded all previous definitions utilized by LoRA. We aim to undertake a comprehensive reconstruction of the entire framework from scratch, ensuring that TSDs are not only well-defined but also effectively utilized in further applications.
3 Build a Framework: Understanding Task-specific Directions
3.1 Preliminaries
Considering a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ where $n<m$, it can be decomposed using SVD as $\mathbf{A}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\mathsf{T}}$. Here, $\mathbf{\Sigma}={\rm diag}(\sigma_{1},\dots,\sigma_{n})$ is the diagonal matrix of singular values, and $\mathbf{U}$ and $\mathbf{V}$ are the left and right singular vectors, respectively. This decomposition can be further expressed as
Since $\{\mathbf{u}_{i}|i=1,2\dots,n\}$ forms the orthogonal bases for $\mathbb{R}^{n}$ and $\{\mathbf{v}_{i}|i=1,2\dots,n\}$ for $\mathbb{R}^{m}$, the matrix space $\mathbb{R}^{n\times m}$ can be spanned by the bases $\{\mathbf{u}_{i}\mathbf{v}_{j}^{\mathsf{T}}|i=1,2\dots,n,j=1,2\dots,m\}$, which are also orthogonal. Thus, matrix $\mathbf{A}$ can be seen as a matrix in a subspace of the original $\mathbb{R}^{n\times m}$ space, spanned by a set of linearly independent bases $\{\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}|i=1,2\dots,n\}$.
Definition 1.
For a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ ($n<m$) with its left and right singular vectors represented by matrices $\mathbf{U}$ and $\mathbf{V}$, respectively, the bases of $\mathbf{A}$ are defined as follows:
  • Core Bases: The core bases of the matrix $\mathbf{A}$ are defined as $\{\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}|i=1,2\dots,n\}$, where each $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}$ is a rank-one matrix formed by the outer product of singular vectors $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$.
  • Global Bases: The global bases of the matrix $\mathbf{A}$ are defined as $\{\mathbf{u}_{i}\mathbf{v}_{j}^{\mathsf{T}}|i=1,2\dots,n,j=1,2\dots,m\}$, covering all combinations of the left and right singular vectors.
Definition 2.
The direction of a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ ($n<m$) is defined based on its global bases, using an expanded set of its singular values padded with zeros, specifically as $(\sigma_{1},0,\dots,0,\sigma_{2},0,\dots,0,\dots,\sigma_{n},\dots,0)\in\mathbb{R}^{nm}$, i.e., flattened $\mathbf{\Sigma}$ by rows.
3.2 What are Task-specific Directions?
We start with the fact that for any specific task, there exists an optimal matrix $\mathbf{W}^{*}\in\mathbb{R}^{n\times m}$ within the matrix space $\mathbb{R}^{n\times m}$ [17]. For a pretrained matrix $\mathbf{W}$, its optimal alteration for this specific task is defined as $\Delta\mathbf{W}^{*}=\mathbf{W}^{*}-\mathbf{W}$. In PEFT, we only possess the information of $\mathbf{W}$ and the directions of $\mathbf{W}$. Since $\Delta\mathbf{W}^{*}$ and $\mathbf{W}^{*}$ are established on their respective bases, we initially project both $\Delta\mathbf{W}^{*}$ and $\mathbf{W}^{*}$ onto the global bases of $\mathbf{W}$, capturing their directions.
Definition 3.
Figure 1: Left: The change rates of $\mathbf{W}$’s core bases, based on $\mathbf{W}^{*}$, vary significantly across each basis, and the bases with higher change rates tend to be concentrated towards the end. Middle: The change rates for the directions corresponding to the smallest 10 singular values of $\mathbf{W}$. The directions associated with the smallest singular values do not always exhibit the highest change rates. Right: After sorting the change rates from highest to lowest, it is evident that only a few directions have significant change rates, while most exhibit very low change rates. The weights are taken from the 16th layer of LLaMA-7B, and the change rates are scaled (Sec. 8.2 for more details).
Established on the global bases of $\mathbf{W}$, $\mathbf{\Pi}_{\mathbf{W}}(\mathbf{W}^{*})$ represents the direction that $\mathbf{W}$ should evolve. $\mathbf{W}$ itself can only utilize up to $n$ bases (i.e., its core bases), meaning that it can at most alter $n$ values of its direction. Thus, we only focus on the changes related to its core directions (for more details, please refer to Sec. 3.2.2). During the transformation, not all core directions’ coordinate values change to the same extent. Given the diversity of downstream tasks and the variability in $\mathbf{W}^{*}$ for different tasks, certain core directions of $\mathbf{W}$ undergo significant changes while others may experience only minimal alterations. Specifically, the change rate $\delta_{i}$ of the coordinate value corresponding to the $i$-th core direction $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}$ of $\mathbf{W}$ is represented as:
$\delta_{i}$ $=|\frac{\mathbf{\Pi}_{\mathbf{W}}(\mathbf{W}^{*})_{ii}-\mathbf{\Pi}_{\mathbf{W}}(\mathbf{W})_{ii}}{\mathbf{\Pi}_{\mathbf{W}}(\mathbf{W})_{ii}+\epsilon}|=|\frac{\mathbf{u}_{i}^{\mathsf{T}}\mathbf{W}^{*}\mathbf{v}_{i}-\mathbf{u}_{i}^{\mathsf{T}}\mathbf{W}\mathbf{v}_{i}}{\mathbf{u}_{i}^{\mathsf{T}}\mathbf{W}\mathbf{v}_{i}+\epsilon}|$ (3)
$=|\frac{\mathbf{u}_{i}^{\mathsf{T}}\Delta\mathbf{W}^{*}\mathbf{v}_{i}}{\mathbf{u}_{i}^{\mathsf{T}}\mathbf{W}\mathbf{v}_{i}+\epsilon}|=|\frac{\mathbf{\Pi}_{\mathbf{W}}(\Delta\mathbf{W}^{*})_{ii}}{\mathbf{u}_{i}^{\mathsf{T}}\mathbf{W}\mathbf{v}_{i}+\epsilon}|.$
Here, $\epsilon=10^{-6}$ is a constant to prevent singular values from being zero. $\sigma_{i}$ is $\mathbf{W}$’s $i$-th singular value, and $\delta_{i}$ represents the change rate of the coordinate value required in the $i$-th core direction $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}$ for transformation from $\mathbf{W}$ to $\mathbf{\Pi}_{\mathbf{W}}(\mathbf{W}^{*})$, specifically quantifies the extent of adaptation needed for a particular task. We hence define the TSD as:
Definition 4.
There are two questions that may arise regarding the definition of TSD as follows:
3.2.1 Why should TSDs be Established on $\mathbf{W}$?
@@
The most accurate direction for TSD should be the direction of $\Delta\mathbf{W}^{*}$. However, since we do not have access to $\Delta\mathbf{W}^{*}$ or $\mathbf{W}^{*}$ during fine-tuning, we only have information about $\mathbf{W}$. Therefore, we establish TSD within the coordinate system based on the directions of $\mathbf{W}$.
@@
3.2.2 Why Must TSDs be in the Core Bases Rather than Other Global Bases?
@@
Indeed, other global bases might also experience significant coordinate changes. However, we do not consider them for the following reasons. First, TSDs are defined based on the change rates. Since the coordinate values of $\mathbf{W}$ on the other global bases are zero, even a minimal projection of $\Delta\mathbf{W}$ onto these global bases could result in an excessively high change rate, which would be unreasonable. Second, if TSDs were defined based on the absolute magnitude of coordinate changes, consider a scenario where a coordinate value in the core basis and a coordinate value in the global basis change by the same amount. If the original coordinate value in the core basis was significantly larger than the change, the change, even if substantial, could be viewed as a minor perturbation relative to the original value. Therefore, the same magnitude of change does not necessarily imply that the corresponding directions are equally important. For these reasons, we define TSDs based solely on the rate of change and restrict their definition to the core bases.
3.3 What are the Properties of TSDs?
To further explore the properties of TSDs, we fully fine-tune LLaMA-7B [4] on commonsense reasoning tasks, and assume that $\mathbf{W}^{*}$ can be obtained through fully fine-tuning. After acquiring the fully fine-tuned weights $\mathbf{W}^{*}$, we compute $\Delta\mathbf{W}^{*}=\mathbf{W}^{*}-\mathbf{W}$, and based on $\Delta\mathbf{W}^{*}$, we calculate the change rates of core directions of $\mathbf{W}$. The results are shown in Fig. 1, where we can draw several conclusions (Please refer to Sec. 8.2 for more details):
@@ 3.3 What are the Properties of TSDs?
Additionally, it is worth noting that the phenomenon presented in Fig. 1 is not an isolated case. Similar behavior has been observed in other tasks and models. For instance, as shown in Fig. 2, results obtained by training Qwen2.5-7B [46] on the math reasoning task [47] exhibit patterns that are fundamentally consistent with our observations in Fig. 1.
@@
Figure 3: We track the precision and recall of predicted directions every 100 training steps in the query, key and value layers of the LLaMA-7B during LoRA fine-tuning, analyzing how well the continuous updated $\Delta\mathbf{W}$ captures TSDs. Left: We compute the average precision/recall across all query, key, and value layers, showing the model’s ability to retain task-specific knowledge for each training step. Across various rank settings of LoRA, these precision/recall rates consistently exceed 0.75/0.70, indicating that $\Delta\mathbf{W}$ reliably captures and integrates TSD information. Right: For a rank setting of $r=32$, we compute the average precision/recall across all steps for each query, key and value layers, revealing their sensitivity to TSDs. The majority of layers maintain an average precision/recall above 0.75/0.70, showing the robustness to capture TSD information.
3.4 Challenges and Explorations: Using TSDs is Not as Easy as Expected
Despite the apparent challenges, we place our confidence in the potential of $\Delta\mathbf{W}$. We hypothesized that the core directions with the highest change rates predicted by $\Delta\mathbf{W}$ from LoRA are strongly associated with TSDs. To test this hypothesis, we conducted extensive experiments [^2] For implementation details, please refer to Sec. 8.3.. We fine-tune LLaMA-7B using LoRA under three ranks, and for every 100 training steps we record the top 8 core directions of $\mathbf{W}$ which exhibit the highest change rates based on the projection of continuously updated $\Delta\mathbf{W}$ (i.e., the directions related to top 8 largest $\delta_{i}=\mathbf{u}_{i}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{i}/\sigma_{i}$, where $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ and $\sigma_{i}$ are the corresponding left/right singular vectors and values of $\mathbf{W}$). These directions serve as the “TSDs” predicted by LoRA. We denote these directions as “launched TSDs” (LTSDs) for differentiation. We then identify the top 4 or 16 TSDs with the highest change rates by fully finetuning as mentioned in Sec. 3.3. We conduct two experimental setups to explore that:
@@ 3.4 Challenges and Explorations: Using TSDs is Not as Easy as Expected
The purpose of the first experiment is more intuitive, aiming to confirm whether the TSDs with the top highest change rates can indeed be captured by $\Delta\mathbf{W}$ from LoRA. The second experiment, which may seem a bit more obscure, is designed to address concerns that, besides capturing the top-ranked TSDs, the LTSDs identified might include irrelevant directions. Therefore, this second setup allows us to examine whether the LTSDs predominantly fall within the upper echelons of TSDs based on their change rates, providing insight into the overall relevance and precision of the LTSDs identified by LoRA. Specifically, we employ two metrics to quantify the quality of LTSDs:
  • Precision: Measures how many of the 8 LTSDs are correctly included in the top 16 TSDs.
@@
The results presented in Fig. 3 lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSDs: